Category Archives: HRS II

Posts about the High Resolution Spectrograph on the HET

The week of Dec 7

Two weeks ago was the HET Board of Director’s meeting in Penn State. The meeting lasted two days and the Board got status reports from HET operations and each of the instrument teams. The main news is that that LRS2 is expecting to reach first science in the first few months of 2016, HRS2 is going to start commissioning before Summer 2016 and VIRUS units are going to be coming in over the next 9 months. The Board was encouraged by the progress being made and hopes that we can continue the pace. They were also pleased to see the progress being made on HPF and were impressed with the clean room facility tour they were given. No major changes or action items were reported by the Board.

Once the commissioning team returned from Happy Valley, we went right back to work and were able to push along the closure of one of our major metrology loops, the guide probes. I am pleased to report that we are able to guide at any telescope Az for full trajectories with the probes at any position within their range.

The week of March 9

The grating and guard rail around the liquid nitrogen tank have been completed. This gives the contractors safer access to the controls during our filling.

The air and glycol (“Blue Thunder”) hoses arrived and installation may start as early as next week. This will actually shut down the structure for a day or two as these are carefully integrated into the azimuth cable wrap below the telescope.

The HRS instrumentation team came out from Austin and disassembled the HRS and packed the optics up. Some of the optics will just be stored while the coated optics are being shipped out to be recoated for the HRS upgrade. The HRS enclosure has been cleared and once some preliminary electrical work is done the inside of the enclosure will be painted flat black.

Our telescope operator and video production guru has completed a new video showing the installation of the temporary clean room frame. See This clean room will house the corrector when it arrives in late April.

A visiting engineer from Texas A&M was back this week with the doors and sides for the VIRUS enclosure. A lot of work continues to seal, plumb and install electrical components in the enclosures and annex.

In night time commissioning we attempted a rewind test. This is to measure the time it takes from the abort of one 20 minute trajectory and rewind and start a 2nd trajectory. It took 79 seconds to make this transition which is a little longer than we really wish but still within the 80 seconds from abort to guider acquisition. We hope to shave 9 seconds off this time so that we can have 10 seconds to have the metrology and guiders acquire their stars.

The mirror team has been getting ahead of their minimum swap rate by swapping another 5 mirrors this last week. They hope to have all of the older mirrors swapped out by the time we start re-commissioning the telescope.

The week of January 26

Happy New Year! The blogger is back from Winter break and travels to January Science meetings.

We will start with the bad news. The tests on the corrector have determined that there is significant asymmetries. These are asymmetries beyond what were found and reported to the board of directors in the December meeting. Those aberrations were going to be removed by changing the plate glass that was going in to seal the bottom of the corrector into a fifth optical element with some small power. That optic has been fabricated and installed in the corrector. The newly discovered astigmatism was found off axis and might have an impact on image quality at the edge of the field of view. To further investigate this optical issue we are adding a few months to the delivery date of the corrector. Instead of being delivered in late January the new delivery date is expected to be early May and significant on-sky commissioning during the rainy season perhaps pushing us into September.

On a more positive note the 2nd VIRUS enclosure has arrived and installed.

Panoramic views showing the VIRUS Enclosures being installed.  The top shows a view from in-front of the mirror and the bottom shows a view from behind the mirror.  Both are taken from the catwalk.

Panoramic views showing the VIRUS Enclosures being installed. The top shows a view from in-front of the mirror and the bottom shows a view from behind the mirror. Both are taken from the catwalk.

There is a lot of plumbing and electrical work required to finish the installation. This is our highest priority.

During the long break the Remote Thermal Area contractors have installed the glycol chiller and are in the process of commissioning it with glycol. This nearly completes the Remote Thermal Area project. Only a few punch list items remain.

The HPF doors also arrived and have been installed. The HPF is now thermally isolated from the spectrograph room. The Penn State Team will monitor the temperatures inside and outside of their enclosure to see if our new Mitsubishi units can hold the temperature to a tight enough tolerance for their specifications. Meanwhile the insulating panels for the HRS II have arrived and will be installed in the coming month. These panels along with an active feed back heater inside the enclosure should allow the HRS II to be held to extremely tight tolerances, < 0.1 C.

The mirror team has started mirror swaps and started using the Strip and Wash room to remove old mirror coatings and prep them for the coating chamber. Our goal is to get the team up to 4 mirrors swaps every 3 weeks.

The week of Oct 27

The contractors continue to work through the Remote Thermal Area punch list including
repairing sheet rock, cleaning up building penetrations, clean up material and reinstalling the roof gutter down spouts. We also have contractors out working on the balance of the in-line fans for the strip and wash duct and the K-hut duct.

We received another shipment of VIRUS Platforms/Ladders;
it contained the platform side supports for both VIRUS right and VIRUS left and a top work platform for one of the VIRUS enclosures. We have been doing test fits of these enclosures in preparation for installing all of them in two weeks.

John Good, mechanical engineer from Austin, was here for the last two weeks to work with the laser tracker. The first days were spent confirming the current mount model then making refinements to it. We now believe we have the best mount model that can be obtained with this laser tracker.

Some work at night was done this week. We went on sky to test the mount model, see if the position of the Celestron Alignment Telescope (CAT) was good enough for future testing. We found the RA, DEC and rho offsets are now working and that the center of rotation falls at the corner of the CAT. Tests of the geosynchronous satellite tracking software within TCS suggests that it is probably working although the CAT is just to small to see these 11-13th mag objects. We believe we are sufficiently ready for any future metrology tests. With this alignment out of the way we put the CCAS tower mirror alignment system back together and have confirmed that all of the components are well aligned and working with alignment to better than 0.5″.

Work in the spectrograph room continues. One of our TOs finished the touch-ups required after the contractors scratched the paint in the area where the HPF enclosure will go. With that done we are ready for the HPF enclosure to arrive next week. We also received the mounts for the HRS sensors. These will be part of the system that monitors and controls the HRS temperature to improve performance in velocity stability.

The really big news is that we installed the large liquid nitrogen tank behind the K-hut. This required days of preparation and coordination between the HET, Crane services, Praxair and Midwest Cryo teams. The two cranes made quick work of putting the crane in place and by lunch the tank was bolted down. In the two following days nearly all of the plumbing to the K-hut vaporizer and the filling location were installed.